(5,2)9(8,3)17(7,6)..... please explain how to get the next number

Answer by Edwin McCravy(20060)

Apparently there is an expression in x and y,
say Ax+By, such that when the point (x,y)=(5,2) 
is substituted, the result is 9, and when the
point (x,y)=(8,3) is substituted, the result is
17. The problem is asking what do we get when 
the point (x,y)=(7,6) is substituted in it?

If that's what is meant, then we substitute 
(x,y)=(5,2) in Ax+By and set that equal to 9:

A(5)+B(2) = 9
5A+2B = 9

Then we substitute (x,y)=(8,3) in Ax+By and set
that equal to 17:

A(8)+B(3) = 17
8A+3B = 17

Then we solve the system:

5A+2B = 9
8A+3B = 17

and get A = 7 and B = -13.

So the expression Ax+By becomes

7x-13y

So we substitute the point (x,y)=(7,6)
in that expression and get:

7(7)-13(6) = 49-78 = -29

Edwin